Freq of autosomal recesive lethal gene is 0.4.find d freq of carrier in a population of 200 individuals.....

Grade:12

1 Answers

Arun
6 years ago
Dear student
 
The frequency of the recessive allele is q = 0.4 (given)
Hence, p = 1-0.4 = 0.6  (as p = 1- q)
According to Hardy- Weinberg principle,  p2 + ?2pq + q2 = 1.
hence,
p2 = (0.6)2 => 0.36
q2 = (0.4)2 => 0.16
2pq = 0.48
since, 2pq is the frequency of hetrozygous carrier,
= 0.48 X 200
=> 96 Individuals

Regards
Arun (askIITians forum expert)